Smith and Wesson are the only two residents in a small neighbourhood. They are considering hiring a security guard. The monthly value of the security guard is $300 to Smith, and $100 to Wesson. Regardless of how the guard is paid, they will protect the entire neighbourhood.

a) Explain what is the maximum price the guard could charge per month and still be hired by one of the neighbours.

b) Suppose the competitive market wage for a security guard is $320 per month. The local government proposes a plan where Smith and Wesson each pay half the monthly fee. Would this plan be voted in? Would the economic surplus be higher if the neighbourhood had a security guard?
Answer #1

a)

Maximum willingness of society = 300 + 100 = 400

Maximum price the guard could charge per month and still be hired by one of the neighbours = 300.

Explanation: Smith is more willing an can pay up to 300.

b).

With payment of half fee (320 / 2), we have:

Net payoff to Smith = 300 - 160 = 140

Net payoff to Wesson = 100 - 160 = - 60 (Loss).

Since Wesson is facing loss, this plan can not be voted in.

Economic surplus with security guard = 300 + 100 - 320 = 80

Yes, economic surplus would be higher by 80 with presence of security guard.

If you are satisfied with the answer, please provide a positive rating. Feel free to comment in case of queries.

Have a nice day ahead!